# WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting

### v2

@@ IV The WRF-GS Framework
In this section, we present the WRF-GS framework for wireless channel modeling based on WRF reconstruction using 3D-GS. This novel framework harnesses the power of 3D-GS and adapts it from the optical domain to the RF domain. The key insight is to convert the space particles into virtual TXs using a continuous volumetric scene function in 3D-GS.
@@ IV-A Problem Description and Framework Overview
We consider a scenario comprising a TX, an RX equipped with an antenna array, and various obstacles such as floors, walls, and furniture. The RX is placed at a fixed position, while the TX is randomly positioned in the space. Following [6], we assume that the distribution of main obstacles in the space is fixed, and any temporary perturbations caused by moving small obstacles are mitigated through techniques like Kalman filtering to minimize their impact. As described in Section III-A, the received signal is the superposition of sub-signals from multiple propagation paths. Each sub-signal can be represented as an LoS signal from a virtual TX. Our objective is to reconstruct the spatial power spectrum received at the RX antenna array by synthesizing the distribution of these virtual TXs within the scenario.
Figure 4: An overview of the WRF-GS framework. The 3D point clouds, which can be randomly distributed or captured via LiDAR sensors, and the position of the TX are first passed into a scenario representation network. This network represents the virtual TXs in the scene using a set of 3D Gaussians, each of which carries environmental attenuation information and signal characteristics. To project these 3D Gaussian representations onto the perception plane of the RX antenna array, the Mercator projection is employed. Finally, the resulting spatial spectra are rendered using the electromagnetic splatting method.
@@
This task is analogous to the optical 3D-GS technique that explicitly models the distribution of particles in a scene to enable arbitrary view synthesis [10]. This motivates us to employ 3D-GS to reconstruct the WRF and propose a novel framework named WRF-GS. Specifically, WRF-GS aims to model the virtual TX distribution to synthesize the received spatial spectrum by combining the EM wave physics and 3D-GS. Since 3D Gaussians possess attributes such as volume, color, and opacity, and serve as the fundamental building blocks in the representation space, they can be naturally translated into virtual TXs in the RF domain. During EM propagation, the overall outcome is perceived as the combined effect of the signal and attenuation, as described in Eqn. (2). In this context, the color attribute is replaced by the signal strength, while the opacity corresponds to the attenuation.
@@
Nevertheless, the projection and splatting of WRF reconstruction in 3D-GS are distinct from the optical domain described in Section III-B. In optical 3D-GS, the goal is to project 3D Gaussian particles representing objects onto the camera’s 2D image plane. This involves applying a coordinate projection transformation that accounts for the camera’s position, orientation, and intrinsic parameters. In contrast, the RF domain requires mapping the virtual TXs represented by 3D Gaussians onto a perception plane of the RX antenna array. Since the RX antenna array receives signals in a hemispherical plane, as discussed in Section III-A, a new projection approach is required to map the 3D Gaussians onto a hemispherical plane, rather than a flat 2D plane. This operation is crucial for aligning the virtual TXs with the specific direction of the RX antenna array. Furthermore, in optical 3D-GS, the rendered image is generated by the splatted Gaussians, including volume, color, and opacity. While the 3D Gaussian attributes in the RF domain represent the signal strength and attenuation. This necessitates a new splatting process to blend the contributions of all the virtual TXs, accounting for their signal strengths and attenuation, to synthesize the received spatial power spectrum.
@@
Our proposed WRF-GS effectively addresses the above challenges and enables the reconstruction of the spatial spectrum. The overall structure of WRF-GS is illustrated in Fig. 4. It takes the positions of physical TXs and 3D point clouds as input and outputs the spatial spectrum at the RX antenna array. WRF-GS consists of the following three core modules:
@@
  • Scenario Representation Network: This module is responsible for representing the virtual TXs using 3D Gaussians. By processing the input data, the 3D Gaussians capture the signal strength and attenuation properties, which enables the synthesis of the WRF in the environment.
@@
  • Projection Model: This model projects the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. To account for the hemispherical nature of the antenna’s reception, we employ the Mercator projection to achieve this spatial projection.
@@
  • Electromagnetic Splatting: Based on the properties of 3D Gaussians and the characteristics of EM wave propagation in space, we employ a hardware acceleration algorithm to efficiently process and combine the contributions of the virtual TXs onto the RX antenna array, ultimately synthesizing the received spatial spectrum.
@@ IV-B Scenario Representation Network
The scenario representation network plays a crucial role in modeling the underlying relationships between the environment and radio signals by representing the virtual TXs as 3D Gaussians using two MLPs. The network architecture, depicted in Fig. 5, follows the DeepSDF structure [22].
To capture the attenuation, which solely depends on the spatial location, we initially feed the positions of the initial point clouds (equivalent to the center coordinates of the 3D Gaussians $G(\boldsymbol{x})$) into a network consisting of eight fully connected layers. Each layer comprises ReLU activations and 128 channels. The output of this network provides the attenuation $\delta(\boldsymbol{x})$ of each input 3D Gaussain $G(\boldsymbol{x})$ at the corresponding position $\boldsymbol{x}$ and a feature vector. Subsequently, we pass this feature vector, along with the TX position $P_{\text{TX}}$, through another two-layer fully connected network (with ReLU activations and 128 and 64 channels, respectively). The final output is the signal $S(\boldsymbol{x})$ that represents the 3D Gaussian signal associated with the virtual TX at position $\boldsymbol{x}$. Therefore, for each 3D Gaussian $G(\boldsymbol{x})$ representing a virtual TX, we obtain the corresponding signal $S(\boldsymbol{x})$ and attenuation $\delta(\boldsymbol{x})$ from the scenario representation network, i.e.,
@@
where $\Theta$ denotes the learnable neural network weights. Considering that signals and attenuation in space are typically expressed as complex numbers, as shown in Eqn. (1), the two outputs here are denoted by $\delta(\boldsymbol{x})=\Delta A(\boldsymbol{x})e^{j\Delta\psi(\boldsymbol{x})}$ and $S(\boldsymbol{x})=A(\boldsymbol{x})e^{j\psi(\boldsymbol{x})}$, respectively, where $A(\boldsymbol{x})$ represents the amplitude and $\psi(\boldsymbol{x})$ represents the phase.
This network is specifically designed to transform the 3D point clouds and TX position information into the signal and attenuation properties of 3D Gaussians, respectively. Unlike DeepSDF, which is typically applied to optics and assumes a fixed TX (light source) position, our network is designed to handle mobile TXs. Moreover, the outputs are all complex-valued signals that incorporate additional phase information. This unique feature enables our model to successfully represent both the signal characteristics of virtual TXs in space and the attenuation information of the surrounding environment.
@@
IV-C Projection Model
The projection model maps the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. There are several key differences compared with the camera model in optical 3D-GS. Unlike pinhole or fisheye cameras, the antenna in our WRF-GS receives signals from a hemispherical direction. In addition, we do not need to know the distortion caused by coordinate transformation. This is because we only care about the signal at the corresponding angular resolution and ignore the signal between angles. Moreover, as shown in Eqn. (4), the arrangement of pixels in the final spatial spectrum differs from conventional photographs that adhere to human visual perception.
To project a 3D point $\boldsymbol{t}={[{t}_{x},{t}_{y},{t}_{z}]}^{T}$ onto the perception plane of the RX antenna array at $\boldsymbol{p}={[{p}_{x},{p}_{y}]}^{T}$, we employ the Mercator projection, a widely used orthographic cylindrical projection in cartography. This allows us to map the latitude and longitude grid of a sphere onto a cylindrical surface, which can then be unrolled onto a flat plane. We define a right-handed Cartesian coordinate at the receiving end of the antenna, with the antenna array oriented in the positive Z-direction and the signal being received in a spherical plane, as shown in Fig. LABEL:Proj(a). Using longitude $\Omega_{\text{lon}}$ and latitude $\Omega_{\text{lat}}$, we can concisely and accurately represent positional relationships in space, which is expressed using inverse trigonometric functions as:
@@ IV-C Projection Model
where ${t}_{r}=\sqrt{{t}_{x}^{2}+{t}_{y}^{2}+{t}_{z}^{2}}$ denotes the distance from the center coordinate of the 3D Gaussian in space to the center of the antenna coordinate, $\arctan 2(\cdot)$ is the 4-quadrant inverse tangent. Since we only consider signals from the upper hemisphere, we have $-\pi\leq\Omega_{\text{lon}}<\pi$ and $20\leq\Omega_{\text{lat}}<\pi/$ 2, and the lower hemisphere is ignored in the following coordinate transformations. Subsequently, we convert the latitude-longitude coordinates (see Fig. LABEL:Proj(b)) into uniform coordinates (see Fig. LABEL:Proj(c)) as
@@
such that $-1\leq{s}_{x}<1$ and $10\leq{s}_{y}<$ 1. The final step of the projection is to transform the uniform coordinates into pixel coordinates of arbitrary resolution (see Fig. LABEL:Proj(d)) as
@@
where $W$ and $H$ represent the number of pixels contained in the width and height of the image, respectively. Considering a one-degree angular resolution, we set $W=360$ and $H=90$.
Through the above coordinate transformations, we maintain the relationship between the signal angles, ensuring that each pixel value in the spatial spectrum corresponds to the AoA of the signal. This allows us to calculate the signals arriving in each direction at the RX antenna array in the next subsection.
@@ IV-D Electromagnetic Splatting
After the coordinate projection transformation, the 3D Gaussians representing the virtual TXs are mapped onto the 2D plane. Note that each 2D Gaussian may cover multiple pixels. In other words, the value of each pixel is the result of the combined contributions from multiple 2D Gaussians. Therefore, the role of electromagnetic splatting module is to efficiently compute the value of each pixel in the 2D plane, which is then utilized to reconstruct the spatial spectrum.
As illustrated in Fig. 9, electromagnetic splatting incorporates the EM wave propagation into the differentiable rasterisation algorithm [10]. To facilitate fast parallel computing, we group adjacent pixels into non-overlapping “tiles” that can be computed independently, as depicted in Fig. 9(a). For each 2D Gaussian, we first determine which tiles it covers, duplicate the Gaussian, and assign the copies to the corresponding tiles. These 2D Gaussians are then sorted by depth order within each tile, as shown in Fig. 9(b). In the scenario representation network, we obtain the signal $S(\boldsymbol{x}_{i})$ and signal attenuation $\delta(\boldsymbol{x}_{i})$ of each 3D Gaussian $G(\boldsymbol{x}_{i})$ at an arbitrary position $\boldsymbol{x}_{i}$. Assuming there are a total of $N$ 3D Gaussians contributing to a particular angle, we can express the effect of the $i^{th}$ Gaussian $G(\boldsymbol{x}_{i})$ on the RX by combining Eqn. (2) and Eqn. (7) as
@@
Consequently, for all 3D Gaussians associated with an angle (pixel) $k$ , the combined effect is expressed as
@@
To facilitate a more intuitive understanding of this process, we provide an illustrative example in Fig. 9(c). For the pixel value $R_{1}$ in the upper-left corner, only the signal $S(\boldsymbol{x}_{1})$ from $G(\boldsymbol{x}_{1})$ contributes, without any attenuation. For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered. The signal from $G(\boldsymbol{x}_{1})$ has no attenuation, but the signals from $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are attenuated by $\delta(\boldsymbol{x}_{1})$ and $\delta(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{2})$, respectively, due to the depth ordering. The computation for the remaining pixel values $R_{k}$ follows a similar principle.

### v4

@@ IV The WRF-GS Framework
In this section, we present the WRF-GS framework, a novel approach for wireless channel modeling based on the reconstruction of WRF using 3D-GS. This framework leverages the power of 3D-GS to explicitly characterize the spatial distribution of the WRF in a given environment. The key innovation lies in transforming spatial particles into virtual TXs using a continuous volumetric scene representation enabled by 3D-GS.
Figure 4: An overview of the WRF-GS framework. The 3D points, which can be randomly distributed or captured via LiDAR sensors, and the position of the TX are first passed into a scenario representation network. This network represents the virtual TXs in the scene using a set of 3D Gaussians, each of which carries environmental attenuation information and signal characteristics. To project these 3D Gaussian representations onto the perception plane of the RX antenna array, the Mercator projection is employed. Finally, the resulting spatial spectra are rendered using the electromagnetic splatting method.
@@ IV-A Problem Description and Framework Overview
We consider a wireless communication scenario comprising a TX, an RX equipped with an antenna array, and various obstacles such as floors, walls, and furniture. The RX is positioned at a fixed location, while the TX is randomly placed within the environment. Following [7], we assume that the primary obstacles in the environment remain stationary, and any transient perturbations caused by moving small obstacles are mitigated using techniques such as Kalman filtering to minimize their impact. As described in Section III-A, the received signal at the RX is a superposition of sub-signals from multiple propagation paths. Each sub-signal can be represented as a LoS signal originating from a virtual TX. Our goal is to reconstruct the spatial power spectrum received at the RX antenna array by modeling the distribution of these virtual TXs within the environment.
@@
This task is analogous to the optical 3D-GS technique, which explicitly models the distribution of particles in a scene to enable arbitrary view synthesis [11]. This motivates us to adapt 3D-GS for WRF reconstruction and propose the novel WRF-GS framework. Specifically, WRF-GS models the distribution of virtual TXs to synthesize the received spatial spectrum by integrating the EM wave physics with 3D-GS. Since 3D Gaussians possess attributes such as volume, color, and opacity, serving as the fundamental building blocks in the representation space, they can be naturally reinterpreted as virtual TXs in the RF domain. During EM propagation, the overall signal is perceived as the combined effect of signal strength and attenuation, as described in Eqn. (2). In this context, the color attribute of a Gaussian corresponds to the signal strength, while the opacity represents the attenuation.
@@
However, the projection and splatting processes for WRF reconstruction in 3D-GS differ significantly from those in the optical domain, as described in Section III-B. In optical 3D-GS, the goal is to project 3D Gaussian particles representing objects onto a 2D image plane, which involves a coordinate projection transformation based on the camera’s position, orientation, and intrinsic parameters. In contrast, the RF domain requires mapping the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. Since the RX antenna array receives signals in a hemispherical plane, as discussed in Section III-A, a new projection approach is needed to map 3D Gaussians onto a hemispherical plane rather than a flat 2D plane. This operation is critical for aligning the virtual TXs with the specific directional sensitivity of the RX antenna array. In addition, while optical 3D-GS renders images by blending the volume, color, and opacity of splatted Gaussians, the RF domain requires blending the signal strengths and attenuations of virtual TXs to synthesize the received spatial power spectrum. This necessitates a new splatting process tailored to the properties of EM wave propagation.
@@
Our proposed WRF-GS framework effectively addresses the aforementioned challenges and enables the accurate reconstruction of the spatial spectrum. The overall architecture of WRF-GS is illustrated in Fig. 4. It takes as input the positions of physical TXs and a set of randomly initialized 3D points representing spatial particles within the environment, and outputs the spatial spectrum at the RX antenna array. WRF-GS consists of the following three core modules:
@@
  • Scenario Representation Network: This module represents the virtual TXs using 3D Gaussians. By processing the input data, the 3D Gaussians capture the signal strength and attenuation properties, enabling the synthesis of the WRF in the environment.
@@
  • Projection Model: This model projects the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. To account for the hemispherical nature of the antenna’s reception, we employ the Mercator projection to achieve this spatial mapping.
@@
  • Electromagnetic Splatting: Leveraging the properties of 3D Gaussians and the characteristics of EM wave propagation, this module employs a hardware-accelerated algorithm to efficiently combine the contributions of the virtual TXs onto the RX antenna array, ultimately synthesizing the received spatial spectrum.
@@ IV-B Scenario Representation Network
The scenario representation network plays an important role in modeling the intricate relationships between the environment and radio signals by representing virtual TXs as 3D Gaussians using two MLPs. The network architecture, depicted in Fig. 5, is inspired by the DeepSDF structure [26] but tailored for wireless channel modeling.
The network begins by processing the initial 3D points, which represent randomly sampled spatial particles within the environment. These points serve as the center coordinates of the 3D Gaussians $G(\boldsymbol{x})$ at location $\boldsymbol{x}$ and are used to initialize the Gaussian primitives. These positions are fed into a neural network to capture the attenuation properties, which depend solely on the spatial location. To characterize these properties, we introduce a new attenuation attribute $\delta(\boldsymbol{x})$ for each 3D Gaussian, which quantifies the signal attenuation caused by environmental interactions at position $\boldsymbol{x}$. This neural network consists of eight fully connected layers, each with ReLU activations and 128 channels. The output of this neural network includes the attenuation $\delta(\boldsymbol{x})$ of each input 3D Gaussain $G(\boldsymbol{x})$ and a feature vector that captures environmental information. Next, this feature vector, along with the TX position $P_{\text{TX}}$, is passed through a second two-layer fully connected network. The first layer has 128 channels and the second layer has 64 channels, both employing ReLU activations. The final output is the signal $S(\boldsymbol{x})$, which represents the 3D Gaussian signal associated with the virtual TX at position $\boldsymbol{x}$. Therefore, for each 3D Gaussian $G(\boldsymbol{x})$ representing a virtual TX, the scenario representation network produces the corresponding signal $S(\boldsymbol{x})$ and attenuation $\delta(\boldsymbol{x})$, as expressed by
@@
where $\Theta$ denotes the learnable weights of the network. Since signals and attenuation in space are typically expressed as complex numbers, as shown in Eqn. (1), the two outputs are expressed as $\delta(\boldsymbol{x})=\Delta A(\boldsymbol{x})e^{j\Delta\psi(\boldsymbol{x})}$ and $S(\boldsymbol{x})=A(\boldsymbol{x})e^{j\psi(\boldsymbol{x})}$, respectively, where $A(\boldsymbol{x})$ represents the amplitude and $\psi(\boldsymbol{x})$ represents the phase.
This network is specifically designed to transform the 3D points (representing spatial particles) and TX position information into the signal and attenuation properties of 3D Gaussians. Unlike DeepSDF, which is typically applied in optic domains and assumes a fixed TX (light source) position, our network is uniquely adapted to handle mobile TXs. Moreover, the outputs are complex-valued signals that incorporate phase information, a critical feature for accurately representing wireless propagation. This capability enables our model to successfully capture both the signal characteristics of virtual TXs and the attenuation properties of the surrounding environment, making it well-suited for wireless channel modeling.
IV-C Projection Model
The projection model is responsible for mapping the virtual TXs represented by 3D Gaussians onto the perception plane of the RX antenna array. There are several key differences compared with the camera model used in optical 3D-GS. First, unlike pinhole or fisheye cameras, the RX antenna in our WRF-GS receives signals from a hemispherical direction. Second, we do not account for distortions caused by coordinate transformations, as we are only concerned with the signal at specific angular resolutions and ignore signals between discrete angles. Third, as shown in Eqn. (4), the arrangement of pixels in the final spatial spectrum differs from conventional photographs, which are designed for human visual perception. Instead, the spatial spectrum is structured to represent signal strength as a function of azimuth and elevation angles.
@@ IV-C Projection Model
To project a 3D point $\boldsymbol{t}={[{t}_{x},{t}_{y},{t}_{z}]}^{T}$ onto the perception plane of the RX antenna array at $\boldsymbol{p}={[{p}_{x},{p}_{y}]}^{T}$, we employ the Mercator projection, a widely used orthographic cylindrical projection in cartography [27]. This allows us to map the latitude and longitude grid of a sphere onto a cylindrical surface, which can then be unrolled onto a flat plane. We define a right-handed Cartesian coordinate system at the receiving end of the antenna, with the antenna array oriented in the positive Z-direction and signals received in a spherical plane, as shown in Fig. LABEL:Proj(a). Using longitude $\Omega_{\text{lon}}$ and latitude $\Omega_{\text{lat}}$, we can concisely and accurately represent positional relationships in space. By applying inverse trigonometric functions, we obtain
@@
where ${t}_{r}=\sqrt{{t}_{x}^{2}+{t}_{y}^{2}+{t}_{z}^{2}}$ denotes the distance from the center coordinate of the 3D Gaussian in space to the origin of the antenna coordinate system. In addition, the function $\arctan 2(\cdot)$ is the 4-quadrant inverse tangent. Since we only consider signals from the upper hemisphere, the longitude and latitude ranges are constrained to $-\pi\leq\Omega_{\text{lon}}<\pi$ and $20\leq\Omega_{\text{lat}}<\pi/$ 2. Signals from the lower hemisphere are ignored in subsequent coordinate transformations. Next, we convert the latitude-longitude coordinates (see Fig. LABEL:Proj(b)) into uniform coordinates (see Fig. LABEL:Proj(c)) as follows:
@@
such that $-1\leq{s}_{x}<1$ and $10\leq{s}_{y}<$ 1. The final step of the projection is to transform the uniform coordinates into pixel coordinates of arbitrary resolution (see Fig. LABEL:Proj(d)), i.e.,
@@
where $W$ and $H$ represent the number of pixels contained in the width and height of the image, respectively. For a one-degree angular resolution, we set $W=360$ and $H=90$.
Through the above coordinate transformations, we maintain the relationship between the signal angles, ensuring that each pixel value in the spatial spectrum corresponds to the AoA of the signal. This allows us to calculate the signal strength arriving from each direction at the RX antenna array, as detailed in the following subsection.
@@ IV-D Electromagnetic Splatting
After the coordinate projection transformation, the 3D Gaussians representing the virtual TXs are mapped onto a 2D plane. Note that each 2D Gaussian may cover multiple pixels. In other words, the value of each pixel is determined by the combined contributions of multiple overlapping 2D Gaussians. The role of the electromagnetic splatting module is to efficiently compute the value of each pixel on the 2D plane, which is then used to reconstruct the spatial spectrum.
As illustrated in Fig. 9, electromagnetic splatting incorporates the principles of EM wave propagation into a differentiable rasterisation algorithm [11]. To enable fast parallel computing, we group adjacent pixels into non-overlapping “tiles”, which are processed independently, as depicted in Fig. 9(a). For each 2D Gaussian, we first identify the tiles it covers, duplicate the Gaussian, and assign these copies to the corresponding tiles. The 2D Gaussians within each tile are then sorted by depth, as shown in Fig. 9(b).
In the scenario representation network, we obtain the signal $S(\boldsymbol{x}_{i})$ and signal attenuation $\delta(\boldsymbol{x}_{i})$ of each 3D Gaussian $G(\boldsymbol{x}_{i})$ at an arbitrary position $\boldsymbol{x}_{i}$. Assuming there are a total of $N$ 3D Gaussians contributing to a specific angle (pixel), we can express the effect of the $i^{th}$ Gaussian $G(\boldsymbol{x}_{i})$ on the RX by combining Eqn. (2) and Eqn. (7) as
@@
Consequently, for all 3D Gaussians associated with a specific angle (pixel) $k$ , the combined effect is expressed as
@@
To facilitate a more intuitive understanding of this process, we provide an illustrative example in Fig. 9(c). For the pixel value $R_{1}$ in the upper-left corner, only the signal $S(\boldsymbol{x}_{1})$ from $G(\boldsymbol{x}_{1})$ contributes, without any attenuation. For the pixel value $R_{2}$ in the upper-right corner, the contributions from $G(\boldsymbol{x}_{1})$, $G(\boldsymbol{x}_{2})$, and $G(\boldsymbol{x}_{3})$ must be considered. The signal from $G(\boldsymbol{x}_{1})$ experiences no attenuation, while the signals from $G(\boldsymbol{x}_{2})$ and $G(\boldsymbol{x}_{3})$ are attenuated by $\delta(\boldsymbol{x}_{1})$ and $\delta(\boldsymbol{x}_{1})\delta(\boldsymbol{x}_{2})$, respectively, due to the depth ordering. The computation for the remaining pixel values $R_{k}$ follows the same principle.
In simple terms, we consider each projected Gaussian as a virtual TX that can affect multiple angles, while signals from multiple virtual TXs are received at each angle. Depending on their distance to the RX, these signals undergo varying degrees of attenuation, ultimately determining the received signal strength at each angle. As explained in Section III-A, the reconstructed spatial spectrum can be represented as a spatial power distribution function, where the value is proportional to the square of the amplitude of the received signal. By deriving the received signal for each angle using Eqn. (12) and Eqn. (13), we obtain the spatial spectrum described in Eqn. (4), which serves as a representation of the underlying WRF.
IV-E Summary
The WRF-GS framework integrates physical models and deep learning techniques to enable comprehensive environment-aware channel modeling and high-fidelity spatial spectrum synthesis. At the core of this approach lies the establishment of a connection between virtual TXs and 3D Gaussians. The scenario representation network models the spatial distribution of signal propagation and attenuation using 3D Gaussians based on input data. To transform these 3D spatial representations onto a 2D plane, we employ the Mercator map projection, ensuring accurate spatial mapping. Finally, the electromagnetic splatting module employs a parallel and efficient algorithm to compute pixel values for each direction, enabling rapid synthesis of the spatial spectrum.
V The WRF-GS+ Framework
The proposed WRF-GS framework demonstrates remarkable effectiveness in reconstructing the WRF and synthesizing spatial spectra for wireless channel modeling, as shown in Section VI and Section VII. However, WRF-GS faces limitations in accurately modeling high-frequency signal variations, as the static nature of the 3D Gaussians in WRF-GS restricts its ability to capture rapid changes in the channel caused by complex multipath effects. To address this challenge, we propose WRF-GS+, an enhanced framework that introduces deformable 3D Gaussians to better model dynamic signal variations and incorporates electromagnetic splatting with $\alpha$-blending to simplify the modeling process.
V-A Scenario Representation Network Using Deformable 3D Gaussians
While WRF-GS effectively models wireless environments, it struggles to accurately capture the rapid signal variations caused by complex multipath effects. To address this limitation, we introduce deformable 3D Gaussians, which enables the framework to decompose the signal into static components and dynamic components. This decomposition allows the network to separately model large-scale fading (e.g., path loss) and small-scale fading (e.g., multipath effects), significantly improving its ability to handle high-frequency signal changes.
As illustrated in Fig. 10, the scenario representation network initializes Gaussian primitives from 3D points, each assigned with attributes such as opacity, signal strength, rotation, and scaling. Unlike WRF-GS, which introduces an additional attribute to explicitly model signal attenuation, WRF-GS+ directly leverages the opacity attribute to capture attenuation characteristics. The signal strength, rotation, and scaling attributes are determined exclusively by the center position of the 3D Gaussian. Notably, these static components of the signal, corresponding to large-scale fading, depend solely on the relative position of the virtual TX to the RX, as the environmental distribution remains invariant. These attributes are thus used to characterize the static components: signal strength is represented by the original color attribute, while rotation and scaling define the size and orientation of Gaussians, thereby affecting the coverage area of the virtual TXs.